# EnScale: Temporally-consistent multivariate generative downscaling via proper scoring rules

### v2

@@ 4.1.3 Analogues
Analogue-based downscaling [Boe2023AValue] is a statistical-dynamical downscaling method which resamples the existing RCM data. It is based on a nearest neighbor search in the GCM data, and stochastic downscaling is approximated by selecting five similar GCM days and using their corresponding RCM days as probabilistic predictions. The main strength of the analogue approach is that, by construction, it yields samples with the correct spatial pattern, as they are all drawn from the existing RCM data. However, it only reproduces the training data and can never create truly new random fields, and the predicted conditional distributions are only discrete. Finally, due to a lack of perfect “near neighbors”, the analogue benchmark is expected to have a larger prediction error between target RCM data and predicted samples than the true RCM’s conditional distribution. Details on our algorithm and potential limitations are given in Sec. B.7.
@@ 4.1.4 Generative Adversarial Network
Generative Adversarial Networks (GANs) have successfully been used to downscale precipitation fields [Harris2022AForecasts, Leinonen2020, PriceStephanRasp2022IncreasingModels]. As an additional benchmark, we consider the approach presented in [Harris2022AForecasts], which is a Wasserstein GAN with gradient norm penalty (WGAN-GP). In the generator objective function, it includes an additional content loss term, more specifically, the mean-squared error of the mean prediction (over eight generated samples per input field) and the ground truth high-resolution output. We adjust [Harris2022AForecasts] [^1] Code: https://github.com/ljharris23/public-downscaling-cgan to our multivariate setting. As for EnScale, we use five input fields (tas, pr, sfcWind, rsds, psl) as the conditioning inputs to the WGAN-GP. Additionally, we include the one-hot-encoded index of the considered GCM-RCM pair and concatenate it to the five input fields. The outputs are the four target fields (tas, pr, sfcWind, rsds). Similar to [Harris2022AForecasts], we scale the precipitation and surface wind values logarithmically with $\log(1+x)$. In addition, we use min-max normalization to ensure that all variables have values on similar scales. For implementation reasons, we confine the input domain of the GCM inputs to $11.25^{\circ}E$ to $26.25^{\circ}E$ and $31.25^{\circ}N$ to $68.75^{\circ}N$, i.e., a smaller extent than covered in EnScale. We use the same hyperparameters as in [Harris2022AForecasts] with a larger batch size of 16 and train until convergence, reached after 2.4 million training samples.
@@ 4.1.5 Corrective Diffusion Model
As another state-of-the-art ML downscaling approach, we consider CorrDiff, which was shown to perform skillfully in multivariate downscaling [Mardani2025ResidualDownscaling]. The model consists of a UNet to predict the (deterministic) conditional mean $E[Y|X]$ and a diffusion model to predict the residual variation around the mean. We adopt the CorrDiff implementation provided in NVIDIA’s PhysicsNeMo library [^2] Code: https://github.com/NVIDIA/physicsnemo/tree/main/examples/weather/corrdiff for the same setting as in the previous GAN experiment, i.e., five input fields concatenated with the one-hot-econded GCM-RCM pair indices, four output fields, and the confined GCM-input-domain.
As in [Mardani2025ResidualDownscaling], we standardize all variables to z-scores and use the same hyperparameter settings. In particular, we use a base embedding size of 128 for the UNet, channels are multiplied with the factors $[1,2,2,2,2]$, and the attention resolution is defined as 8 (as our input is of size $128\times 128$). We use a smaller batch size of 32, and train until convergence: first, the UNet for 4 million training samples and, afterwards, the diffusion model for 26 million training samples.
@@ 4.2 Validation
To design our validation metrics, we again leverage the connection between generative models and forecast evaluation (Sec. 2.5.1). Two important properties of the generated samples by the ML model are calibration and sharpness [Gneiting2007ProbabilisticSharpness].
@@ 4.2.1 Energy score and CRPS
We consider the multivariate EnScale-loss for the full spatial field for each variable separately (see Sec. 2.5.3). In addition, we calculate CRPS values, which coincide with the energy score in the univariate case. We calculate the CRPS for each variable and in each location and then average it across the spatial field.
@@ 4.2.2 Spatial structure
In addition, we compute power spectral densities (PSD). We follow [Harris2022AForecasts] and consider the radially-average PSD, calculated with pysteps. Following [Harris2022AForecasts], we compare the sampled PSDs with the RCM’s PSDs via the log spectral distance, called RALSD:
@@
Here, $\overline{P}_{\text{RCM, i}}$ and $\overline{P}_{\text{gen, i}}$ are the power spectra for the RCM data and one generated field at one wavelength $i$, where the 2D spectra were averaged radially. $N$ is the number of wavelengths considered. The scaling by factor 10 is chosen such that the resulting units are in decibel. To arrive at RALSD, we compute PSDs separately for each test day and compute the RALSD on daily PSDs as in Eq. (3). Finally, we average over all days in the test set, as in [Harris2022AForecasts].
However, the PSDs $\overline{P}_{\text{gen, i}}$ for different random samples can vary substantially even on the same day, due to the large variability in the conditional $p_{Y|X}$. Thus, RALSD is also sensitive to general mismatches between samples and target, and even a method with correct spatial structure might score poorly if its realizations deviate from the RCM target. Hence, we consider averaging the radially-averaged PSDs across all days before compute the log spectral density. This version emphasizes systematic errors in spatial structure, and we refer to this metric as RALSD-avg.
@@ 4.2.4 Extremes
4.2.5 Temporal structure
We evaluate the temporal structure of the samples by computing auto-correlations. For each variable and each location, we calculate temporal lag-1 autocorrelation for the generated time series as well as the RCM target, and report absolute errors of the samples.
@@ 4.2.6 Variable dependencies
In addition, we consider a novel metric of calibration: we compute RHs and MCB metrics (see Sec. 4.2.3) for the difference of two variables in the samples. However, as directly comparing variables with different physical units is not meaningful, we first apply a data transformation to standardize them. Specifically, for each variable and each dataset (RCM or sampled), we transform the data to a uniform distribution. This ensures that the marginal distributions for each variable are identical, eliminating differences to units. In addition, it removes mismatches between RCM and sampled data in individual variable distributions, such that the remaining discrepancies arise only from different dependence structures between the variables. After transformation, we calculate the differences between two variables. We compute RHs and MCBs for each grid point separately and then averaged across locations.
@@ 4.3 Contributions to variability in EnScale
We aim to better understand the variability EnScale’s samples and present decompositions of its conditional variance $\operatorname{Var}(Y|X)$. As discussed (Sec. 2.1), there is a wide range of possible realizations of high-resolution fields given a particular GCM input. EnScale allows us to quantify this variability. In this section, all variances and expected values refer to those of EnScale’s samples, assuming already that EnScale emulates the true RCM’s distributions accurately. In addition, $\operatorname{Var}$ and $E$ denote quantities for the pooled distribution across multiple GCM-RCM pairs, $\iota$ denotes the index for a GCM-RCM pair, and $\operatorname{Var}^{\iota}$ resp. $E^{\iota}$ refer to distributions by an individual pair.
@@
We aim to decompose the internal variability of EnScale’s full GCM-RCM path into the variability on coarse and local scales. For this, we split the variance of EnScale, $\operatorname{Var}^{\iota}(Y|X)$, into contributions from (i) variance of the coarse model, $\operatorname{Var}^{\iota}(Z|X)$, which we call coarse-scale variability, and (ii) variance of the super-resolution model, $\operatorname{Var}^{\iota}(Y|Z)$, which we call local-scale variability (see Sec. 2.2). As $Z$ and $Y$ have different dimensions and physical meanings, $\operatorname{Var}^{\iota}(Z|X)$ and $\operatorname{Var}^{\iota}(Y|Z)$ are not directly comparable. Instead, we can formally decompose the conditional variance $\operatorname{Var}^{\iota}(Y|X)$ as follows:
@@
where the first term is the variance of $Y|Z$, averaged over all possible $Z$ (conditional on $X$). The second term describes the variance of the conditional mean $E^{\iota}[Y|Z]$, again for different possible $Z$. This is conceptually very similar to $\operatorname{Var}^{\iota}(Z|X)$, except it is “evaluated in YY”. A proof is given in App. B.8. This decomposition holds separately in each grid point and for each variable. We average each of the terms over all grid points as well as over all $X$ in the test set.
We calculate the quantities above as follows: for a given GCM input $X$, we generate 9 realizations of $Z$ with the coarse model. For each $Z$, we again generate 9 realizations of $Y$ with the super-resolution model, and compute $\operatorname{Var}^{\iota}(Y|Z)$ and $\operatorname{Var}^{\iota}(E^{\iota}[Y|Z]|X)$ for each grid point. Afterwards, we average across $X$, grid points and GCM-RCM pairs. To restrict memory usage, the calculation is only done for the first year in the test set.
@@
Thus, the total variance can be written as an average of the variances of the individual RCMs (internal variability) plus the squared deviations of the means of each RCM from the mean of the pooled data (model differences). This decomposition analogous to classical ANOVA. For a proof, see [Casella2024StatisticalInference]. As above, we consider this decomposition separately per grid point and per variable, and average over grid points and predictors $X$.
@@ 5.1 Comparative overview of method performances
Figure 5: Summary of performance of EnScale compared to the benchmarks in several selected categories, shown for the interpolation test period (2030-39). Energy score (see Sec. 5.4.1), Calibration (see Sec. 5.4.3), Spatial structure (see Sec. 5.4.2), Temporal structure (see Sec. 5.4.5), Extremes (see Sec. 5.4.4), Multivariate dependencies (Sec. 5.5). The chosen metrics for the categories are outlined in more detail in the main text. All metrics are normalized such that EnScale attains a score of 1 in the respective category and all other scores are expressed relative to EnScale. In all categories, lower values indicate better performance. As calibration for deterministic models is not meaningful, we do not show it for NN-det.
Figure 6: Time series examples. Solid lines show the RCM time series (green) and a single randomly chosen realization of EnScale (blue), EnScale-t (yellow), and CorrDiff (orange). Shaded areas indicate the minimum and maximum from 9 samples of each model. The figure shows spatially averaged temperature for the first year in the interpolation test set (2030) and for the first GCM-RCM pair (CNRM-CM5 and ALADIN63).
@@
Fig. 5 summarizes the performance of EnScale compared to the benchmarks for several important categories. “Energy score” refers to the spatially multivariate energy score, “Spatial structure” is calculated from the max-pooled CRPS and the radially-average log spectral density (RALSD and RALSD-avg), “Temporal structure” refers to the autocorrelation functions, and “Extremes” contains the marginal quantiles and quantile losses. To evaluate variable dependencies, we compute correlations for multiple variable pairs and assess calibration of their differences. The results are then averaged and summarized as “correlations” and “calibration”. All metrics are analyzed in more detail in the respective sections below. Unless otherwise stated, all metrics are shown on the interpolation test set (2030-2039) and averaged across all GCM-RCM pairs.
@@
We find that EnScale and EnScale-t reach very good scores for many categories and variables in comparison to the benchmarks. It particularly stands out with respect to calibration, extremes as well as the multivariate dependencies where it outperforms benchmarks by a large margin.
@@
CorrDiff yields mixed results: despite good energy score and realistic spatial structure, its performance for calibration, extremes and multivariate dependencies is poor. We find that these deficiencies are a consequence of the sampling scheme in CorrDiff, which is intended to provide ensemble members. The current implementation yields highly autocorrelated time series and restricts the diversity of samples. Thus, we adapted the sampling to allow a broader exploration of the space of possible high-resolution fields. This new version, which we call CorrDiff-s, substantially improves on CorrDiff’s performance in most metrics, as shown in App. A.2.
@@
The analogues benchmark performs well in many metrics as is to be expected. This benchmark resamples existing fields. For this reason, it perfectly reproduces the spatial characteristics and variable correlations of the RCM, thus reaching excellent scores in these categories. The GAN is inferior compared to the other ML-approaches. EasyUQ performs well on metrics that consider each location separately, but leads to the worst performance regarding calibration and spatial structure.
@@
Fidelity regarding the temporal structure is of particular relevance and the extension to EnScale-t exhibits the best scores of all stochastic methods in that regard. This is because EnScale-t is the only method that explicitly models temporal structure. All other methods downscale each day independently of previous days, with CorrDiff being one exception due to shared random initializations between different time steps. Generated time series also show that EnScale-t and CorrDiff produce smoother, and thus more realistic, trajectories than EnScale (see Fig. 6).
We compare EnScale (EnScale-t) to EnScale-no-label (EnScale-t-no-label) which exclude the one-hot-encoded index of the GCM and RCM in the super-resolution model, thus learning a joint super-resolution map for all pairs (see Fig. A.1). EnScale yields small improvements to EnScale-no-label. Overall, results are very similar, again underlining that the super-resolution map behaves similarly for multiple RCMs.
In addition, we repeat all metric calculations on the extrapolation test period (2090-2099) and show the result in App. A.4. As this is only a mild extrapolation case (as training data is until 2089), scores are very similar to the interpolation period and the ordering of the methods stays consistent.
5.2 Computational cost
Method Training Inference
EnScale 24 h 1.75 h
CorrDiff 216 h 36 h
Table 2: Training and inference times for EnScale and CorrDiff. All runtimes were checked on a single NVIDIA A100 GPU (80GB) and exclude data loading. Inference times refer to producing a high-resolution ensemble of 100 years for a single GCM-RCM pair with 10 samples per day.
EnScale aims to be a computationally lightweight alternative to state-of-the-art generative models. As an example, we compare train and inference times of EnScale to the best-performing and most recent ML-baseline, CorrDiff (Tab. 2). We find that training EnScale only takes about 1 day, whereas the diffusion model requires almost 10 times as long. The difference for inference is even larger, with EnScale being around 20 times faster. However, it should be noted that inference times can likely be improved for both models. For EnScale, the sampling can be done in a few minutes, but the current bottleneck is the inversion of the data-transformation.
EnScale is parameter-efficient due to its sparse layers (Sec. 2.4). Despite modeling effects specific to each location, the network only has 66 million parameters. The UNet in our CorrDiff implementation has around 80 million parameters (as in [Mardani2025ResidualDownscaling]) and is used in both the regression as well as the residual modeling step, yielding a total of around 160 million parameters.
Computational efficiency gains for EnScale are due to both the overall setup and the model architecture. The sparse local layers were designed for progressive resolution refinement in a stochastic manner, effective in combination with the energy score loss and our multi-step setup (Sec. 2.2), and are key to achieving the short training time. For example, training a fully convolutional architecture for the last steps of the super-resolution model took several days.
5.3 Exemplary samples
Figure 7: Exemplary samples from EnScale for all variables. The first column presents the unseen target RCM data by ALADIN63 driven by CNRM-CM5 on day 2035-05-06. Columns 2-4 show three corresponding random samples from EnScale. We chose a day with average performance, i.e., the EnScale-loss is roughly equal to the median score of all days in the interpolation test set.
Qualitatively, EnScale outputs realistic spatial fields that resemble the RCM data as illustrated in Fig. 7 for an exemplary GCM input. Examples for two other GCM inputs are shown in Fig. A.4 and A.5. For example, Fig. A.4 demonstrates a case where samples by EnScale are particularly similar to the RCM data (i.e., small EnScale-loss) .
Per GCM-RCM pair, only one RCM field per GCM input is available. Thus, it is not possible to compare the variability in EnScale outputs with the raw data directly. Instead, one can follow the nearest-neighbor approach as in our analogue benchmark. Exemplary samples of this benchmark are shown in Fig. A.6. In this case, the variability stems partly from the small training dataset and imperfect near-neighbors (App. B.7). Nevertheless, the examples give a first impression of the variability that is present in the raw GCM-RCM data.
We present exemplary samples for the other machine learning methods (GAN and CorrDiff) in Fig. A.7 and A.8. We find similar conclusions as for EnScale: while the benchmark’s spatial fields resemble the RCM, they all exhibit considerable variability. GAN outputs tend to be blurry, whereas CorrDiff samples capture fine spatial details more realistically.

### v3

@@ 4.1.3 Analogues
Analogue-based downscaling [9] is a statistical-dynamical downscaling method which resamples the existing RCM data. It is based on a nearest neighbor search in the GCM data, and stochastic downscaling is approximated by selecting five similar GCM days and using their corresponding RCM days as probabilistic predictions. The main strength of the analogue approach is that, by construction, it yields samples with the correct spatial pattern, as they are all drawn from the existing RCM data. However, it only reproduces the training data and can never create truly new random fields, and the predicted conditional distributions are only discrete. Finally, due to a lack of perfect “near neighbors”, the analogue benchmark is expected to have a larger prediction error between target RCM data and predicted samples than the true RCM’s conditional distribution. Details on our algorithm and potential limitations are given in Sec. C.7.
@@ 4.1.4 Generative Adversarial Network
Generative Adversarial Networks (GANs) have successfully been used to downscale precipitation fields [25, 37, 48]. As an additional benchmark, we consider the approach presented in [25], which is a Wasserstein GAN with gradient norm penalty (WGAN-GP). In the generator objective function, it includes an additional content loss term, more specifically, the mean-squared error of the mean prediction (over eight generated samples per input field) and the ground truth high-resolution output. We adjust [25] [^1] Code: https://github.com/ljharris23/public-downscaling-cgan to our multivariate setting. As for EnScale, we use five input fields (tas, pr, sfcWind, rsds, psl) as the conditioning inputs to the WGAN-GP. Additionally, we include the one-hot-encoded index of the considered GCM-RCM pair and concatenate it to the five input fields. The outputs are the four target fields (tas, pr, sfcWind, rsds). Similar to [25], we scale the precipitation and surface wind values logarithmically with $\log(1+x)$. In addition, we use min-max normalization to ensure that all variables have values on similar scales. For implementation reasons, we confine the input domain of the GCM inputs to $11.25^{\circ}E$ to $26.25^{\circ}E$ and $31.25^{\circ}N$ to $68.75^{\circ}N$, i.e., a smaller extent than covered in EnScale. We use the same hyperparameters as in [25] with a larger batch size of 16 and train until convergence, reached after 2.4 million training samples.
@@ 4.1.5 Corrective Diffusion Model
As another state-of-the-art ML downscaling approach, we consider CorrDiff, which was shown to perform skillfully in multivariate downscaling [44]. The model consists of a UNet to predict the (deterministic) conditional mean $E[Y|X]$ and a diffusion model to predict the residual variation around the mean. We adopt the CorrDiff implementation provided in NVIDIA’s PhysicsNeMo library [^2] Code: https://github.com/NVIDIA/physicsnemo/tree/main/examples/weather/corrdiff, commit 469d6aa09354ec91ebe1f388a8247c96e502e9c8 for the same setting as in the previous GAN experiment, i.e., five input fields concatenated with the one-hot-econded GCM-RCM pair indices, four output fields, and the confined GCM-input-domain.
As in [44], we standardize all variables to z-scores and use the same hyperparameter settings. In particular, we use a base embedding size of 128 for the UNet, channels are multiplied with the factors $[1,2,2,2,2]$, and the attention resolution is defined as 8 (as our input is of size $128\times 128$). We use a smaller batch size of 32, and train until convergence: first, the UNet for 4 million training samples and, afterwards, the diffusion model for 26 million training samples.
In it’s original implementation, CorrDiff generates each ensemble member by drawing a single random noise realization that is shared across all time points. This construction implicitly imposes temporal dependence in the generated time series, but variability across realizations is limited. For this reason, we also tested an adapted version which we call CorrDiff-s. In this variant, we randomize the sampling scheme: For each day, noise is drawn independently of previous days and of other ensemble members. This version represents the variance of samples more faithfully. However, it removes temporal autocorrelation, leading to negative errors in autocorrelations, similar to the other benchmarks.
@@ 4.2 Validation
To design our validation metrics, we again leverage the connection between generative models and forecast evaluation (Sec. 2.5.1). Two important properties of the generated samples by the ML model are calibration and sharpness [20].
@@ 4.2.1 Energy score and CRPS
We consider the multivariate EnScale-loss for the full spatial field for each variable separately (see Sec. C.2), i.e., treating all grid cells as a multivariate vector. In addition, we calculate CRPS values, which coincide with the energy score in the univariate case. We calculate the CRPS for each variable and in each location and then average it across the spatial field.
@@ 4.2.2 Spatial structure
In addition, we compute power spectral densities (PSD). We follow [25] and consider the radially-average PSD, calculated with pysteps. Following [25], we compare the sampled PSDs with the RCM’s PSDs via the log spectral distance, called RALSD:
@@
Here, $\overline{P}_{\text{RCM, i}}$ and $\overline{P}_{\text{gen, i}}$ are the power spectra for the RCM data and one generated field at one wavelength $i$, where the 2D spectra were averaged radially. $N$ is the number of wavelengths considered. The scaling by factor 10 is chosen such that the resulting units are in decibel.
We compute PSDs separately for each test day and consider two options for averaging PSDs and spectral distances: (i) For the RALSD-metric, we compute the RALSD on daily PSDs as in Eq. (2). Finally, we average over all days in the test set, as in [25]. (ii) Alternatively, we average radially-averaged PSDs across all days, plot these and compute the log spectral density of the averaged spectra. This distance emphasizes systematic errors in spatial structure, and we refer to this metric as RALSD-avg.
PSDs $\overline{P}_{\text{gen, i}}$ for different random samples can vary substantially even on the same day, due to the large variability in the conditional $p_{Y|X}$. Thus, RALSD is also sensitive to deviations from individual samples from the target, whereas RALSD-avg focuses more directly on systematic differences in the average spatial structure.
@@ 4.2.3 Calibration
Alternatively, ranks could be pooled across spatial locations. However, opposite biases at different locations could cancel out in the pooled histogram, thus making the result difficult to interpret. Hence, we instead compute the MCB for each location individually, before averaging this across locations.
@@ 4.2.5 Variable dependencies
In addition, we consider a novel metric of calibration: we compute RHs and MCB metrics (see Sec. 4.2.3) for the difference of two variables in the samples. However, as directly comparing variables with different physical units is not meaningful, we first apply a data transformation to standardize them. Specifically, for each variable and each dataset (RCM or sampled), we transform the data to a uniform distribution via the ECDFs (similar to the above data transformation). This ensures that the marginal distributions for each variable are identical, eliminating differences to units. In addition, it removes mismatches between RCM and sampled data in individual variable distributions, such that the remaining discrepancies arise only from different dependence structures between the variables. After transformation, we calculate the differences between two variables. We compute RHs and MCBs for each grid point separately and then averaged across locations.
4.2.6 Temporal structure
We evaluate the temporal structure of the samples by computing auto-correlations. For each variable and each location, we calculate temporal autocorrelation functions (ACFs) for the generated time series as well as the RCM target for lags up to 20 days. We consider both raw as well as high-pass filtered time series. For the latter, we subtract a running mean over 30 days first and calculate ACFs on the filtered time series. We report both signed as well as absolute errors of the ACFs in the samples compared to the RCM target.
@@ 4.3 Contributions to variability in EnScale
We aim to better understand the variability EnScale’s samples and present decompositions of its conditional variance $\operatorname{Var}(Y|X)$, with $Y$ being the high-resolution RCM data and $X$ the GCM input, as above. As discussed (Sec. 2.1), there is a wide range of possible realizations of high-resolution fields given a particular GCM input. EnScale allows us to quantify this variability. In this section, all variances and expected values refer to those of EnScale’s samples, assuming already that EnScale emulates the true RCM’s distributions accurately. In addition, $\operatorname{Var}$ and $E$ denote quantities for the pooled distribution across multiple GCM-RCM pairs, $\iota$ denotes the index for an RCM, quantities with a superscript $\iota$, such as $\operatorname{Var}^{\iota}$ and $E^{\iota}$, refer to distributions by an individual RCM.
@@
We aim to decompose the internal variability of EnScale’s full GCM-RCM path into the variability on coarse and local scales. Consider an input $X$ from one of the GCMs. Recall that $Z$ denotes the coarsened RCM data. For the input $X$, we split the variance of EnScale, $\operatorname{Var}^{\iota}(Y|X)$, into contributions from (i) variance of the coarse model, $\operatorname{Var}^{\iota}(Z|X)$, which we call coarse-scale variability, and (ii) variance of the super-resolution model, $\operatorname{Var}^{\iota}(Y|Z)$, which we call local-scale variability (see Sec. 2.2). As $Z$ and $Y$ have different dimensions and physical meanings, $\operatorname{Var}^{\iota}(Z|X)$ and $\operatorname{Var}^{\iota}(Y|Z)$ are not directly comparable. Instead, we can formally decompose the conditional variance $\operatorname{Var}^{\iota}(Y|X)$ as follows:
@@
where the first term is the variance of $Y|Z$, averaged over all possible $Z$ (conditional on $X$). The second term describes the variance of the conditional mean $E^{\iota}[Y|Z]$, again for different possible $Z$. This is conceptually very similar to $\operatorname{Var}^{\iota}(Z|X)$, except it is “evaluated in YY”. A proof is given in App. C.8. This decomposition holds separately in each grid point and for each variable. We average each of the terms over all grid points as well as over all $X$ in the test set (i.e. over all days, but also over all GCM runs in our dataset).
We calculate the quantities above as follows: for a given GCM input $X$, we generate 9 realizations of $Z$ with the coarse model. For each $Z$, we again generate 9 realizations of $Y$ with the super-resolution model, and compute $\operatorname{Var}^{\iota}(Y|Z)$ and $\operatorname{Var}^{\iota}(E^{\iota}[Y|Z]|X)$ for each grid point. Afterwards, we average across $X$, grid points and GCM-RCM pairs. To restrict memory usage, the calculation is only done for the first half of the test set.
@@
Thus, the total variance can be written as an average of the variances of the individual RCMs (internal variability) plus the squared deviations of the means of each RCM from the mean of the pooled data (model differences). This decomposition analogous to classical ANOVA. For a proof, see [10]. As above, we consider this decomposition separately per grid point and per variable, and average over grid points and predictors $X$. In this case, we only report results for one driving GCM (and two corresponding RCM datasets).
@@ 5 Evaluating the performance of EnScale
5.1 Visual evaluation of samples
Figure 5: Samples from EnScale for all variables. The first column presents the unseen target RCM data by ALADIN63 driven by CNRM-CM5 on day 2035-05-06. Columns 2-4 show three corresponding random samples from EnScale. We chose a day with average performance, i.e., the EnScale-loss is roughly equal to the median score of all days in the interpolation test set.
Qualitatively, EnScale outputs realistic spatial fields that resemble the RCM data as illustrated in Fig. 5 for one GCM input. Examples for two other GCM inputs are shown in Fig. A.1 and A.2. For example, Fig. A.1 demonstrates a case where samples by EnScale are particularly similar to the RCM data (i.e., small EnScale-loss).
Per GCM-RCM pair, only one RCM field per GCM input is available. Thus, it is not possible to compare the variability in EnScale outputs with the raw data directly. Instead, one can follow the nearest-neighbor approach as in our analogue benchmark. Samples of this benchmark are shown in Fig. A.3. In this case, the variability stems partly from the small training dataset and imperfect near-neighbors (App. C.7). Nevertheless, the examples give a first impression of the variability that is present in the raw GCM-RCM data.
We present samples for the other machine learning methods (GAN and CorrDiff) in Fig. A.4 and A.5. We find similar conclusions as for EnScale: while the benchmark’s spatial fields resemble the RCM, they all exhibit considerable variability. GAN outputs tend to be blurry, whereas CorrDiff samples capture fine spatial details more realistically.
@@ 5.2 Comparative overview of method performances
Figure 6: Summary of performance of EnScale compared to the benchmarks in several selected categories, shown for the interpolation test period (2030-39). Energy score (see Sec. 5.3.1), Calibration (see Sec. 5.3.4), Spatial structure (see Sec. 5.3.2), Temporal structure (see Sec. 5.5), Extremes (see Sec. 5.3.5), Multivariate dependencies (Sec. 5.4). The chosen metrics for the categories are outlined in more detail in the main text. First, all individual metrics are normalized by the value of EnScale-t, such that EnScale-t attains a score of 1 for each metric, and all other scores are expressed relative to it. The normalized metrics are then averaged in each category. In all categories, lower values indicate better performance. As calibration for deterministic models is not meaningful, we do not show it for NN-det.
@@
Fig. 6 provides an overview of the performance of EnScale compared to the benchmarks for several important categories. Scores are normalized such that EnScale-t receives a 1 in each metric, and all other methods are expressed relative to it. The category “Energy score” refers to the energy score computed over the full spatial field (all grid cells jointly, separately for each variable, Tab. 2). The score for “Spatial structure” is calculated from the max-pooled CRPS (Tab. 2) and the radially-average log spectral density (RALSD and RALSD-avg, Tab. 2 and Tab. A.1), “Temporal structure” refers to the error in autocorrelation (Tab. A.4), and “Extremes” contains the errors in marginal quantiles and miscalibration metrics for quantiles (Tab. A.2). [^3] We removed the error in the 1 % quantile for precipitation, as small absolute errors were blown up to unrealistically large numbers with the relative scaling. To evaluate variable dependencies, we compute correlations for multiple variable pairs and assess calibration of their differences (Tab. A.3). The results are then averaged and summarized as “correlations” and “calibration”. These values are intended as a summary for comparison purposes and do not include formal uncertainty estimates. Standard errors are reported in the metric-specific tables below. All metrics are analyzed in more detail in the respective sections below. Unless otherwise stated, all metrics in this and the following sections are shown on the interpolation test set (2030-2039) and averaged across all GCM-RCM pairs.
@@
We find that EnScale and EnScale-t achieve consistently competitive scores across all categories and variables. In the majority of cases, they rank among the top three methods and beat most benchmarks. They particularly stands out with respect to calibration, extremes, multivariate dependencies. For temporal structure, EnScale-t outperforms all benchmarks by a large margin.
@@
CorrDiff-s often reaches best results except for temporal structure. It beats all other methods for calibration, spatial structure and extremes, and achieves results similar to EnScale for energy scores and multivariate dependencies. CorrDiff yields mixed results: despite good energy score and realistic spatial structure, its performance for calibration, extremes and multivariate dependencies is poor. We find that these deficiencies are a consequence of the sampling scheme in CorrDiff, which is intended to provide ensemble members. Due to the large conditional variance in the distribution of RCM fields given GCM inputs, the original CorrDiff sampling scheme is suboptimal, as it yields highly autocorrelated time series and restricts the diversity of samples. An independent sampling strategy, as in CorrDiff-s, better captures the spread of possible high-resolution fields.
@@
The analogues benchmark performs well in many metrics as is to be expected. This benchmark resamples existing fields. For this reason, it perfectly reproduces the spatial characteristics and variable correlations of the RCM, thus reaching excellent scores in these categories. The GAN is inferior compared to the other ML-approaches. EasyUQ performs well on metrics that consider each location separately, but leads to the worst performance regarding calibration and spatial structure. As expected, NN-det is inferior to the benchmarks for the probabilistic metrics, in particular for the extremes and for the energy score. Also, its spatial structure for the variables with higher variance (all except temperature) is worse than for EnScale.
@@
Fidelity regarding the temporal structure is of particular relevance for some statistics. The extension to EnScale-t exhibits the best scores of all stochastic methods for temporal autocorrelations. This is because EnScale-t is the only method that explicitly models temporal structure. All other methods downscale each day independently of previous days, with CorrDiff being one exception due to shared random initializations between different time steps. Generated time series also show that EnScale-t and CorrDiff produce smoother, and thus more realistic, trajectories than EnScale (see Sec. 5.5) and CorrDiff-s.